# Primitive Quantum Gates for an $SU(3)$ Discrete Subgroup: $\Sigma(36\times 3)$

### v1

@@ I Introduction
To this end, several digitization methods have been proposed in the past decade to render the bosonic Hilbert space finite. Traditionally, most approaches have considered only qubit devices. However, due to recent demonstrations demonstration of qudit gates, there has been an increasing interest in qudit-based digitization methods [9, 10, 11, 12, 13, 14, 15, 16]. One digitization method utilizes the representation (electric field) basis, and impose a cut off on the maximal representation [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 13, 32, 33, 34, 35, 36]. Another recent proposal [37, 16], the $q$-deformed formulation renders a finite dimension by replacing the continuous gauge group with a so-called quantum group [38]. Moreover, the loop-string-hadron formulation explicitly enforces gauge invariance on the Hilbert space [39, 40, 41, 42, 43] before truncation. Each of the above methods is extendable to the full gauge group, i.e. it has an infinite-dimensional limit. Other methods with begin with different formulations or perform different approximations exist such as light-front quantization [44, 45, 46], conformal truncation [47], strong-coupling and large-$N_{c}$ expansions [48, 49].
@@
Another promising approach to digitization is the discrete subgroup approximation [65, 66, 67, 68, 69, 70, 71, 72, 14, 73, 74, 75, 65, 66, 67, 68, 69, 70, 71, 72, 14, 73, 74, 76, 77, 78, 79, 39]. This method was explored early on in Euclidean lattice field theory to reduce computational resources in Monte Carlo simulations of gauge theories. Replacing $U(1)$ by $\mathbb{Z}_{N}$ was considered in [80, 81]. Extensions to the crystal-like subgroups of $SU(N)$ were made in Refs. [82, 83, 84, 66, 67, 69, 85, 86, 87], including with fermions [88, 89]. Theoretical studies revealed that the discrete subgroup approximation corresponds to continuous groups broken by a Higgs mechanism [90, 91, 92, 93, 94]. On the lattice, this causes the discrete subgroup to poorly approximate the continuous group below a freezeout lattice spacing $a_{f}$ (or beyond a coupling $\beta_{f}$), see Fig. 1.
@@
Figure 1: Euclidean calculations of lattice energy density $\langle E_{0}\rangle$ of $\Sigma(36\times 3)$ as measured by the expectation value of the plaquette as a function of Wilson coupling $\beta$ on $8^{d}$ lattices for (top) $(2+1)$-dimensions (bottom) $(3+1)$-dimensions. The shaded region indicates $\beta\geq\beta_{s}$.
@@ II Properties of $\Sigma(36\times 3)$
where $0\leq p,~{}q,~{}r\leq 2$ and $0\leq s,~{}t\leq 1$. This indicates that either 8 qubits (2 each for $p$ , $q$, and $r$, and one each for $s$ and $t$) or 3 qutrits ($p$, $q$, and $r$) and 2 qubits ($s$ and $t$) will be required to store the group register. Because the indices $p,~{}q,$ and $r$ take on values between 0 and 2, there exists an ambiguity in mapping the three-level states to a pair of two-level systems. We use the mapping $|0\rangle_{3}=|00\rangle_{2}$, $|1\rangle_{3}=|01\rangle_{2}$, and $|2\rangle=|10\rangle_{2}$ with the $|11\rangle_{2}$ state being forbidden. Throughout this work, we will use $|\rangle_{3}$ to denote a three-level state and $|\rangle_{2}$ to denote a two-level state when there is the possibility of ambiguity. In this way, the index $p$ is decomposed in binary as $p=p_{0}+2p_{1}$ and encoded as the state $|p\rangle_{3}=|p_{1}p_{0}\rangle_{2}$. This process is done similarly for $q$ and $r$.
@@
$EC=$ $\omega CE,~{}VC=EV,~{}\text{and}~{}VE=C^{2}V.$ (6)
$V^{2s+t}C^{q}=$ $C^{(1+s)q(1-t)}E^{t(1+s)q}V^{2s+t},$
$E^{r}C^{q}=$ $\omega^{rq}C^{q}E^{r},$
$V^{2s+t}E^{r}=$ $C^{t(s+1)(3-r)}E^{(1-t)r(1+s)}V^{2s+t}.$ (7)
Table 2: Character table of $\Sigma(36\times 3)$ with $\omega=e^{2\pi i/3}$ [103].
@@ IV Primitive Gates
The final gate required is the group Fourier transform $\mathfrak{U}_{F}$. The Fourier transform of a finite $G$ is defined as
@@
The second sum is taken over $\rho$, the irreducible representations of $G$; $\hat{f}$ denotes the Fourier transform of $f$. After performing the Fourier transform, the register is denoted as a $\hat{G}$-register to indicate the change of basis. A schematic example of this gate is show in Fig. 2.
@@ IV.1 Inversion Gate
A detailed derivation of the permutation rules and the associated $\mathfrak{U}_{-1}$ is found in App. A along with two other forms of $\mathfrak{U}_{-1}$ which use fewer ancilla. The idealized qubit circuit is shown in Fig. 3 and requires 119 T-gates and 4 clean ancilla. The qubit-qutrit hybrid encoding $\mathfrak{U}_{-1}$ is found in Fig. 4.
@@ IV.2 Multiplication Gate
where $\mathfrak{U}_{\times,O}$ indicate multiplying the state of the $O$ generator register from the $g_{1}$ register onto the $g_{2}$ register. The breakdown of using this method and the product rules from Eq. (IV.2) yields the circuits composed in Fig. 5 and 6 for the two encodings.
Figure 5: Qubit implementation of $\mathfrak{U}_{\times}$ using the permutation gate $\chi$ and its inverse (both shaded orange) using 2 ancilla and has cost of 308 T-gates.
Figure 6: Qutrit+Qubit implementation of $\mathfrak{U}_{\times}$ using the permutation gate $\chi=X^{(0,1)}X^{(1,2)}$ and its inverse (both shaded orange).
@@ IV.3 Trace Gate
Then, the trace operator can be written as $\mathfrak{U}_{Tr}\left(\theta\right)=\exp{\left(-\mathrm{i}\,\theta\,H_{{\rm Tr}}\right)}$. To obtain the matrix form of $\hat{H}_{Tr}$, one may fix a basis $\left\{|g_{1}\rangle,{...},|g_{|G|}\rangle\right\}$ where $|G|=108$ is the order of the group. In this basis, $\hat{H}_{Tr}$ is diagonal, and each diagonal entry is given by $H_{i,i}=Tr(g_{i})$.
To obtain a quantum circuit realizing $\mathfrak{U}_{{\rm Tr}}$, we use the tree-traversal algorithm developed in [115] [^3] The codes accompanying the cited publication can be found on this Github repository. which was shown to yield an exact circuit with asymptotically optimal CNOT gates count. The circuit obtained has 130 CNOT gates and 111 $R_{z}$ gates. Additional methods are also found in Ref. [116]. A second method for deriving this gate involves mapping group elements to their respective trace classes. $\Sigma(36\times 3)$ has 14 conjugacy classes that map to 10 different trace classes. If we only require the real part of the trace then this grouping reduces the 10 trace classes to 7 trace classes. The seven valid traces are ${\rm ReTr}(g)=\{3,-\frac{3}{2},0,\pm 1,\pm\frac{1}{2}\}$, which we can be labelled using three bits $\left(v_{0},v_{1},v_{2}\right)$ as shown in Table 3.
$\operatorname{Re}\operatorname{Tr}(g)$ $-3$ $-\frac{3}{2}$ $-1$ $-\frac{1}{2}$ 0 1 $\frac{1}{2}$
@@
This map can be represented as three boolean functions, one for each of the variables $v_{0}$, $v_{1}$ and $v_{2}$. For quantum computation, it is convenient to write boolean functions in the so-called exclusive-or sum of products (ESOP) form [117, 118]. Then, the function can be mapped to a quantum circuit in a straightforward manner since each term in the ESOPs corresponds to a Toffoli gate. For each function, we start with their minterm forms [119]. Then, we use the exorcism algorithm to find a simpler ESOP expression for each of the three functions. After factorizations, we show the final expressions in Eq. 26.
@@
Figure 7 shows the quantum circuit of the operator $V$ realizing the map $\left(p,q,r,s,t\right)\mapsto\left(v_{0},v_{1},v_{2}\right)$. Finally, the circuit of $\mathfrak{U_{Tr}}^{\prime}(\theta)$ is shown in Figure 8.
@@ IV.4 Fourier Transform Gate
Since $\Sigma(36\times 3)$ has 108 elements, on a qubit device $\mathfrak{U}_{F}$ must be embedded into a larger $2^{8}\times 2^{8}=256\times 256$ matrix [^4] An explicit construction of the matrix $\mathfrak{U}_{F}$ is found in the supplementary information. While the matrix This matrix was then passed to the Qiskit v0.43.1 transpiler, and an optimized version of $\mathfrak{U}_{F}$ needed 30956 CNOTs, 2666 $R_{X}$, 32806 $R_{Y}$, and 55234 $R_{Z}$ gates; the Fourier gate is the most expensive qubit primitive. As will be discussed in Sec. V, $\mathfrak{U}_{F}$ dominates the total simulation costs and future work should be devoted to finding a $\Sigma(36\times 3)$ $\mathfrak{U}_{QFT}$
For the hybrid qubit-qutrit implementation, $\mathfrak{U}_{F}$ is of dimensions $108\times 108$. To obtain a quantum circuit, we built a qubit-qutrit compiler, see Ref. [122]. The outline of the compilation is discussed in Appendix B. The compiler uses the gates discussed in Sec. III, and Table 4 shows its resulting gates count.
Basic Gate $\mathcal{U}_{-1}$ $\mathcal{U}_{\times}$ $\mathcal{U}_{Tr}$ $\mathcal{U}_{F}$
@@
Table 4: Gate cost of primitive gates for $\Sigma(36\times 3)$ for a qutrit-qubit architecture. The costs for $\mathcal{U}_{F}$ were obtained with the hybrid compiler described in Appendix B. $C_{d}X_{b,c}$ refers to an $X$ rotation between states $b,c$ of a qutrit controlled by a qudit with $d$ levels
@@ V Resource Costs
The relatively deep circuits presented above strongly suggest that simulating $\Sigma(36\times 3)$ will require error correction. The preclusion of universal transversal sets of gates stated in the Eastin-Knill theorem [123] requires compromises be made. In most error correcting codes, the Clifford gates are designed to be transversal [104, 124, 125, 126, 127]. This leaves the nontransversal T gate as the dominant cost of fault-tolerant algorithms [104, 128]. Beyond these standard codes, novel universal sets exist with transversal $\mathbb{BT},\mathbb{BO},\mathbb{BI}$ and He(3) [^5] The Heisenberg group of dimension 3, which is a noncrystal-like subgroup of SU(3) gates [129, 130, 131, 132, 133] which warrant exploration for use in lattice gauge theory.
@@
$\mathcal{U}_{-1}$ 119 4
$\mathcal{U}_{\times}$ 308 2
$\mathcal{U}_{Tr}$ $210+8.05\log_{2}(1/\epsilon)$ 7
$\mathcal{U}_{F}$ $185897.5\log_{2}(1/\epsilon)$ 0
For this work we will consider the following decompositions of gates into T gates for our resource estimates. First, while the CNOT is transversal, the Toffoli gate decomposes into six CNOTs and seven T gates [104]. With this, one can construct any CnNOT gates using $12\lceil\log_{2}{n}\rceil-$ 1 Toffoli gates and $n-2$ dirty ancilla qubits [^6] A clean ancilla is in state |$\ket{0}$. Dirty ancillae have unknown states. which can be reused later [108, 104, 109]. For the $R_{Z}$ gates, we use the repeat-until-success method of [134] which finds these gates can be approximated to precision $\epsilon$ with on average $1.15\log_{2}(1/\epsilon))$ T gates (and at worst $-9+4\log_{2}(1/\epsilon)$ [135]). For $R_{Y}$ and $R_{X}$, one can construct them with at most 3 $R_{Z}$. Putting all together, we can construct gate estimates for $\Sigma(36\times 3)$ (See Tab. 5).

### v4

@@ I Introduction
To this end, several digitization methods have been proposed in the past decade to render the bosonic Hilbert space finite. Traditionally, most approaches have considered only qubit devices. However, due to recent demonstrations of qudit gates, there has been an increasing interest in qudit-based digitization methods [9, 10, 11, 12, 13, 14, 15, 16]. One digitization method utilizes the representation (electric field) basis, and impose a cut off on the maximal representation [17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29, 30, 31, 13, 32, 33, 34, 35, 36]. Another recent proposal [37, 16], the $q$-deformed formulation renders a finite dimension by replacing the continuous gauge group with a so-called quantum group [38]. Moreover, the loop-string-hadron formulation explicitly enforces gauge invariance on the Hilbert space [39, 40, 41, 42, 43] before truncation. Each of the above methods is extendable to the full gauge group, i.e. it has an infinite-dimensional limit. Other methods begin with different formulations or perform different approximations exist such as light-front quantization [44, 45, 46], conformal truncation [47], strong-coupling and large-$N_{c}$ expansions [48, 49].
@@
Another promising approach to digitization is the discrete subgroup approximation [65, 66, 67, 68, 69, 70, 71, 72, 14, 73, 74, 75, 65, 66, 67, 68, 69, 70, 71, 72, 14, 73, 74, 76, 77, 78, 79, 39]. This method was explored early on in Euclidean lattice field theory to reduce computational resources in Monte Carlo simulations of gauge theories. Replacing $U(1)$ by $\mathbb{Z}_{N}$ was considered in [80, 81]. Extensions to the crystal-like subgroups of $SU(N)$ were made in Refs. [82, 83, 84, 66, 67, 69, 85, 86, 87], including with fermions [88, 89]. Theoretical studies revealed that the discrete subgroup approximation corresponds to continuous groups broken by a Higgs mechanism [90, 91, 92, 93, 94]. We additionally provide properties of certain $SU(3)$ discrete groups in Tab. 1. On the lattice, this causes the discrete subgroup to poorly approximate the continuous group below a freezeout lattice spacing $a_{f}$ (or beyond a coupling $\beta_{f}$), see Fig. 1.
@@
Figure 1: Euclidean calculations of lattice energy density $\langle E_{0}\rangle$ of $\Sigma(36\times 3)$ as measured by the expectation value of the plaquette as a function of Wilson coupling $\beta$ on $8^{d}$ lattices for (top) $(2+1)$-dimensions (bottom) $(3+1)$-dimensions. The shaded region indicates $\beta\geq\beta_{f}$.
@@ II Properties of $\Sigma(36\times 3)$
where $0\leq p,~{}q,~{}r\leq 2$ and $0\leq s,~{}t\leq 1$. This indicates that either 8 qubits (2 each for $p$ , $q$, and $r$, and one each for $s$ and $t$) or 3 qutrits ($p$, $q$, and $r$) and 2 qubits ($s$ and $t$) will be required to store the group register. Because the indices $p,~{}q,$ and $r$ take on values between 0 and 2, there exists an ambiguity in mapping the three-level states to a pair of two-level systems. We use the mapping $|0\rangle_{3}=|00\rangle_{2}$, $|1\rangle_{3}=|01\rangle_{2}$, and $|2\rangle_{3}=|10\rangle_{2}$ with the $|11\rangle_{2}$ state being forbidden. Throughout this work, we will use $|\rangle_{3}$ to denote a three-level state and $|\rangle_{2}$ to denote a two-level state when there is the possibility of ambiguity. In this way, the index $p$ is decomposed in binary as $p=p_{0}+2p_{1}$ and encoded as the state $|p\rangle_{3}=|p_{1}p_{0}\rangle_{2}$. This process is done similarly for $q$ and $r$.
@@
The Hamiltonians we particularly are targeting are the pure gauge theory Kogut Susskind Hamiltonian,
$H_{KS}=$ $\sum_{\square}Tr(g_{1}g_{2}g_{3}^{\dagger}g_{4}^{\dagger})|g_{1}g_{2}g_{3}g_{4}\rangle\langle g_{1}g_{2}g_{3}g_{4}|$
$+\sum_{l}\sum_{g_{l},h_{l}}e^{\beta Tr(g_{l}h_{l}^{\dagger})}|g_{l}\rangle\langle h_{l}|,$ (8)
where $\sum_{\square}$ indicates a sum over all of the plaquettes with $g_{1},~{}...,g_{4}$ elements of the plaquettes. Additionally, the second term is the kinetic term where the sum over $l$ is a sum over all links. There is generally a freedom in the construction of the electric term. In Appendix D, we provide a straightforward construction based on the procedure outlined in Ref. [104]. The second Hamiltonian we consider is the improved Hamiltonian, $H_{I}$, which was highlighted in Ref. [96]. This includes terms with the six link rectangles and an extended electric field operator. The desire to consider improved Hamiltonians comes from the fact that there are both reduced lattice spacing errors, i.e., the discretization artifacts are moved to $\mathcal{O}(a^{4})$, and the $\beta_{f}$ is a larger value.
Table 2: Character table of $\Sigma(36\times 3)$ with $\omega=e^{2\pi i/3}$ [103]. Size indicates the number of elements in the group while Ord. (order) indicates the number of times the operator can be multiplied before yielding the identity.
@@ IV Primitive Gates
The next gate required is the group Fourier transform $\mathfrak{U}_{F}$. The Fourier transform of a finite $G$ is defined as
@@
The second sum is taken over $\rho$, the irreducible representations of $G$; $\hat{f}$ denotes the Fourier transform of $f$. After performing the Fourier transform, the register is denoted as a $\hat{G}$-register to indicate the change of basis. A schematic example of this gate is show in Fig. 2. A related final gate is $\mathfrak{U}_{Ph}$ which induces the phases corrresponding to the eigenvalue sof the kinetic term of the Hamiltonian.
@@ IV.1 Inversion Gate
A detailed derivation of the permutation rules and the associated $\mathfrak{U}_{-1}$ is found in App. A along with two other forms of $\mathfrak{U}_{-1}$ which use fewer ancilla. The idealized qubit circuit is shown in Fig. 3 and requires 119 T-gates and 4 clean ancilla. [^3] A clean ancilla is in state |$\ket{0}$. Dirty ancillae have unknown states. The qubit-qutrit hybrid encoding $\mathfrak{U}_{-1}$ is found in Fig. 4.
@@ IV.2 Multiplication Gate
where $\mathfrak{U}_{\times,O}$ indicate multiplying the state of the $O$ generator register from the $g_{1}$ register onto the $g_{2}$ register. We provide a detailed discussion of the breakdown of the rules in App. B. The breakdown of using this method and the product rules from Eq. (IV.2) yields the circuits composed in Fig. 5 and 6 for the two encodings.
Figure 5: Qubit implementation of $\mathfrak{U}_{\times}$ using the permutation gate $\chi$ and its inverse (both shaded orange) using 2 ancillae and has a cost of 308 T-gates.
Figure 6: Qutrit+Qubit implementation of $\mathfrak{U}_{\times}$ using the permutation gate $\chi=X^{(1,2)}X^{(0,1)}$ and its inverse (both shaded orange).
@@ IV.3 Trace Gate
Then, the trace operator can be written as $\mathfrak{U}_{Tr}\left(\theta\right)=\exp{\left(-\mathrm{i}\,\theta\,H_{{\rm Tr}}\right)}$. This operator corresponds to the phasing of the magnetic plaquette operator when $g$ corresponds to a closed Wilson loop. To obtain the matrix form of $\hat{H}_{Tr}$, one may fix a basis $\left\{|g_{1}\rangle,{...},|g_{|G|}\rangle\right\}$ where $|G|=108$ is the order of the group. In this basis, $\hat{H}_{Tr}$ is diagonal, and each diagonal entry is given by $H_{i,i}=Tr(g_{i})$.
To obtain a quantum circuit realizing $\mathfrak{U}_{{\rm Tr}}$, we use the tree-traversal algorithm developed in [116] [^4] The codes accompanying the cited publication can be found on this Github repository. which was shown to yield an exact circuit with asymptotically optimal CNOT gates count. The circuit obtained has 130 CNOT gates and 111 $R_{z}$ gates. Additional methods are also found in Ref. [117]. A second method for deriving this gate involves mapping group elements to their respective trace classes. $\Sigma(36\times 3)$ has 14 conjugacy classes that map to 10 different trace classes. If we only require the real part of the trace then this grouping reduces the 10 trace classes to 7 trace classes. The seven valid traces are ${\rm ReTr}(g)=\{3,-\frac{3}{2},0,\pm 1,\pm\frac{1}{2}\}$, which we can be labelled using three bits $\left(v_{0},v_{1},v_{2}\right)$ as shown in Tab. 3.
$\operatorname{Re}\operatorname{Tr}(g)$ $3$ $-\frac{3}{2}$ $-1$ $-\frac{1}{2}$ 0 1 $\frac{1}{2}$
@@
This map can be represented as three boolean functions, one for each of the variables $v_{0}$, $v_{1}$ and $v_{2}$. For quantum computation, it is convenient to write boolean functions in the so-called exclusive-or sum of products (ESOP) form [118, 119]. Then, the function can be mapped to a quantum circuit in a straightforward manner since each term in the ESOPs corresponds to a Toffoli gate. For each function, we start with their minterm forms [120]. Then, we use the exorcism algorithm to find a simpler ESOP expression for each of the three functions. After factorizations, we show the final expressions in the following equation:
@@
Figure 7 shows the quantum circuit of the operator $V$ realizing the map $\left(p,q,r,s,t\right)\mapsto\left(v_{0},v_{1},v_{2}\right)$. Finally, the circuit of $\mathfrak{U_{Tr}}^{\prime}(\theta)$ is shown in Fig. 8.
@@ IV.4 Fourier Transform Gate
Since $\Sigma(36\times 3)$ has 108 elements, on a qubit device $\mathfrak{U}_{F}$ must be embedded into a larger $2^{8}\times 2^{8}=256\times 256$ matrix [^5] An explicit construction of the matrix $\mathfrak{U}_{F}$ is found in the supplementary information. While the matrix was then passed to the Qiskit v0.43.1 transpiler, and an optimized version of $\mathfrak{U}_{F}$ needed 30956 CNOTs, 2666 $R_{X}$, 32806 $R_{Y}$, and 55234 $R_{Z}$ gates; the Fourier gate is the most expensive qubit primitive. As will be discussed in Sec. V, $\mathfrak{U}_{F}$ dominates the total simulation costs and future work should be devoted to finding a $\Sigma(36\times 3)$ $\mathfrak{U}_{QFT}$
For the hybrid qubit-qutrit implementation, $\mathfrak{U}_{F}$ is of dimensions $108\times 108$. To obtain a quantum circuit, we built a qubit-qutrit compiler, see Ref. [123]. The outline of the compilation is discussed in Appendix E. The compiler uses the gates discussed in Sec. III, and Table 4 shows its resulting gates count. The final component is a phasing corresponding to the gauge kinetic term. This involves decomposing the diagonal phasing operation into a sum of Pauli matrices or equivalents for qudit based systems. For the pure qubit based system this decomposition involves 256 $R_{z}$ rotations and 254 CNOT gates. The gate cost on a mixed qubit-qutrit device is shown in Table 8.
Basic Gate $\mathfrak{U}_{-1}$ $\mathcal{U}_{\times}$ $\mathcal{U}_{Tr}$ $\mathcal{U}_{F}$
@@
Table 4: Gate cost of primitive gates for $\Sigma(36\times 3)$ for a qutrit-qubit architecture. The costs for $\mathcal{U}_{F}$ were obtained with the hybrid compiler described in Appendix E. $C_{d}X_{b,c}$ refers to an $X$ rotation between states $b,c$ of a qutrit controlled by a qudit with $d$ levels
@@ V Resource Costs
The relatively deep circuits presented above strongly suggest that simulating $\Sigma(36\times 3)$ will require error correction and longer coherence times on quantum devices. The preclusion of universal transversal sets of gates stated in the Eastin-Knill theorem [124] requires compromises be made. In most error correcting codes, the Clifford gates are designed to be transversal [105, 125, 126, 127, 128]. This leaves the nontransversal T gate as the dominant cost of fault-tolerant algorithms [105, 129]. Beyond these standard codes, novel universal sets exist with transversal $\mathbb{BT},\mathbb{BO},\mathbb{BI}$ and He(3) [^6] The Heisenberg group of dimension 3, which is a noncrystal-like subgroup of SU(3) gates [130, 131, 132, 133, 134] which warrant exploration for use in lattice gauge theory.
@@
$\mathfrak{U}_{-1}$ 119 4
$\mathfrak{U}_{\times}$ 308 2
$\mathfrak{U}_{Tr}$ $378+8.05\log_{2}(1/\epsilon)$ 7
$\mathfrak{U}_{F}$ $185898\log_{2}(1/\epsilon)$ 0
$\mathfrak{U}_{Ph}$ $294.4\log_{2}(1/\epsilon)$ 0
For this work, we will consider the following decompositions of gates into T gates for our resource estimates. First, while the CNOT is transversal, the Toffoli gate decomposes into six CNOTs and seven T gates [105]. With this, one can construct any CnNOT gates using $12\lceil\log_{2}{n}\rceil-$ 1 Toffoli gates and $n-2$ dirty ancilla qubits which can be reused later [109, 105, 110]. For the $R_{Z}$ gates, we use the repeat-until-success method of [135] which finds these gates can be approximated to precision $\epsilon$ with on average $1.15\log_{2}(1/\epsilon))$ T gates (and at worst $-9+4\log_{2}(1/\epsilon)$ [136]). For $R_{Y}$ and $R_{X}$, one can construct them with at most three $R_{Z}$. Putting all together, we can construct gate estimates for $\Sigma(36\times 3)$ (See Tab. 5).